x0.210 1-216 Vo.210 -2.060 m/s (c) t-0.315 s x0.315 = -487 Vo.315-3.090 m/s Need Help? ReadIt 2 P030.W A baseball is hit so that it travels straight upward after being struck by the bat. A fan observes that it takes 3.20 s for the ball to reach its maximum height. (a) Find the balls initial velocity. m/s upward (b) Find the height it reaches. Need Help? ReadIt Submit Answer SaveProgress」 Pr Submit Assignm
Answer #1

(a) Suppose 'u' is the initial velocity of the ball.

Now, at the maximum height final velocity, v of the ball will be equal to ZERO.

time to reach the maximum height, t = 3.20 s

Through-out the upward and downward journey of the ball, gravitational acceleration 'g' will acts on the ball.

Use the expression -

v = u + a*t

=> 0 = u + g*3.20 = u - 9.8*3.20

=> u = 9.8*3.20 = 31.36 m/s upward

(b) Maximum height reached by the ball, h(max) = u^2 / (2*g) = 31.36^2 / (2*9.8) = 50.18 meter (Answer)